OpenStudy (anonymous):

please, help me find answer: integral (sqrt (36-x^2)xdx). I think, there the first step is: integral (sqrt(36-x^2)dx^2. And what will then?

OpenStudy (anonymous):

\[\huge \int\limits \sqrt{36-x^2}xdx\]this it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Haha :) Ever heard of u-substitution? :)

OpenStudy (anonymous):

yes:) but i don't know how it will be there)

OpenStudy (anonymous):

Well, let's play with it a bit, this will help us in the future... \[\huge \int\limits (\sqrt{36-x^2} \ )\left(-\frac{1}{2} \right)(-2x)dx\] Essentially the same, right? :)

OpenStudy (anonymous):

I just multiplied 2 and divided by 2, negative :)

OpenStudy (anonymous):

yes, i understand)

OpenStudy (anonymous):

Now, -1/2 is a constant, we can bring it out, right? :) \[\huge -\frac{1}{2}\int\limits (\sqrt{36-x^2} \ ) (-2x)dx\]

OpenStudy (anonymous):

yes, bring it out)

OpenStudy (anonymous):

Now, can you see where u-substitution comes in? :)

OpenStudy (anonymous):

hm, noo...

OpenStudy (anonymous):

Here's a hint. Let \[\large u = 36-x^2\] Can you continue?

OpenStudy (anonymous):

I will try)

OpenStudy (anonymous):

what will v-substitution?

OpenStudy (anonymous):

u, v, it doesn't really make a difference. I usually use u first, though

OpenStudy (anonymous):

So...\[\large u = 36-x^2\] Now find \[\large\frac{du}{dx}\]

OpenStudy (anonymous):

-2x

OpenStudy (anonymous):

precisely :) \[\large \frac{du}{dx}=-2x\]\[\large du = -2xdx\] Can you do it now? :)

OpenStudy (anonymous):

I understand it) but i don't know about second substitution

OpenStudy (anonymous):

No need for that... back to the integral. \[\huge -\frac{1}{2}\int\limits (\sqrt{36-x^2} \ ) (-2x)dx\] \[\large u = 36-x^2\]\[\large du=(-2x)dx\] Replace what can be replaced :)

OpenStudy (anonymous):

now i will try)

OpenStudy (anonymous):

yees, i could :)

OpenStudy (anonymous):

Awesome :)

OpenStudy (anonymous):

thank you very much!!:)

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

I'm very glad:) i'm newcomer there:)